OpenStudy (ria23):

Find the x-intercepts for the equation y=-x^2+8x-12

geerky42 (geerky42):

x-intercepts is where y=0, so you have \[-x^2+8x-12=0\] Multiply both sides by -1, then try to factor.

OpenStudy (ria23):

\[x ^{2}-8x+12=-1 \] ? ;o

geerky42 (geerky42):

So multiplying 0 by -1 will get you -1?

OpenStudy (ria23):

xD Oh... Oops. . .

geerky42 (geerky42):

Haha, so yeah; \(x^2-8x+12=0\) Can you factor it?

OpenStudy (ria23):

(x-2)(x-6) ?

geerky42 (geerky42):

Yep, so you have \((x-2)(x-6)=0\) So either \((x-2)\) or \((x-6)\)must be equal to 0. So you have two cases: \[x-2=0\\~\\x-6=0\] Can you find the value of x?

OpenStudy (ria23):

Positive 2 and 6...

geerky42 (geerky42):

Yep, that's your answer: (2,0) and (6,0)

geerky42 (geerky42):

Welcome.

OpenStudy (ria23):

^_^ Thank yhu!~